OpenStudy (anonymous):

Find the quotient x^2-100/x+10

OpenStudy (anonymous):

top is difference of 2 squares = (x-10)(x+10) / (x+ 10) = x-10

OpenStudy (anonymous):

IS that the actual answer? Doesn't make sense to me.. ?

OpenStudy (anonymous):

well (x+ 10)'s cancel out leaving x-10 (x+10) / (x+10) = 1

OpenStudy (anonymous):

suzie where are you?